Hi! How to do the bank loan , 1st JAN 2011 the business borrowed 100000 on a 5 year loan.
interest rate of 6% per annum was payable monthly. It was on an interest only basis with interest paid monthly at the end of each month.
Chat started
1
Expert's answer
2020-01-21T07:36:30-0500

Annuity scheme:

1. Calculate the% rate per month: 6% / 12 = 0.005

2. Calculate the monthly payment: 100,000 * 0.005 / 1- (1 + 0.005) -60 = 1933.28

3. We calculate the total amount of payments on the loan = 1933.28 * 60 = 115 993.92

4. Overpayment size: 115,993.92-100,000 = 15,993.92


0 *01.01.2011 1 933.28 1 423.69 509.59 98 576.31 0

1 01.02.2011 1 933.28 1 430.95 502.33 97 145.36 0

2 01.03.2011 1 933.28 1 486.15 447.13 95 659.22 0

3 01.04.2011 1 933.28 1 445.81 487.47 94 213.41 0

4 *01.05.2011 1 933.28 1 468.67 464.61 92 744.74 0

5 01.06.2011 1 933.28 1 460.66 472.62 91 284.08 0

6 01.07.2011 1 933.28 1 483.11 450.17 89 800.96 0

7 01.08.2011 1 933.28 1 475.66 457.62 88 325.30 0

8 01.09.2011 1 933.28 1 483.18 450.10 86 842.12 0

9 *01.10.2011 1 933.28 1 505.02 428.26 85 337.10 0

10 01.11.2011 1 933.28 1 498.41 434.87 83 838.69 0

11 01.12.2011 1 933.28 1 519.83 413.45 82 318.86 0

12 *01.01.2012 1 933.28 1 513.83 419.45 80 805.03 0

13 01.02.2012 1 933.28 1 522.63 410.65 79 282.40 0

14 01.03.2012 1 933.28 1 556.36 376.92 77 726.03 0

15 *01.04.2012 1 933.28 1 538.28 395.00 76 187.75 0

16 01.05.2012 1 933.28 1 558.59 374.69 74 629.17 0

17 01.06.2012 1 933.28 1 554.02 379.26 73 075.15 0

18 *01.07.2012 1 933.28 1 573.89 359.39 71 501.26 0

19 01.08.2012 1 933.28 1 569.91 363.37 69 931.34 0

20 *01.09.2012 1 933.28 1 577.89 355.39 68 353.45 0

21 01.10.2012 1 933.28 1 597.12 336.16 66 756.34 0

22 01.11.2012 1 933.28 1 594.03 339.25 65 162.31 0

23 *01.12.2012 1 933.28 1 612.81 320.47 63 549.50 0

24 01.01.2013 1 933.28 1 610.30 322.99 61 939.21 0

25 01.02.2013 1 933.28 1 617.64 315.64 60 321.56 0

26 01.03.2013 1 933.28 1 655.64 277.64 58 665.92 0

27 01.04.2013 1 933.28 1 634.33 298.96 57 031.60 0

28 01.05.2013 1 933.28 1 652.03 281.25 55 379.57 0

29 *01.06.2013 1 933.28 1 651.07 282.21 53 728.50 0

30 01.07.2013 1 933.28 1 668.32 264.96 52 060.18 0

31 01.08.2013 1 933.28 1 667.99 265.29 50 392.19 0

32 *01.09.2013 1 933.28 1 676.49 256.79 48 715.71 0

33 01.10.2013 1 933.28 1 693.04 240.24 47 022.67 0

34 01.11.2013 1 933.28 1 693.66 239.62 45 329.01 0

35 *01.12.2013 1 933.28 1 709.74 223.54 43 619.27 0

36 01.01.2014 1 933.28 1 711.00 222.28 41 908.27 0

37 *01.02.2014 1 933.28 1 719.72 213.56 40 188.55 0

38 *01.03.2014 1 933.28 1 748.30 184.98 38 440.25 0

39 01.04.2014 1 933.28 1 737.39 195.89 36 702.85 0

40 01.05.2014 1 933.28 1 752.28 181.00 34 950.57 0

41 *01.06.2014 1 933.28 1 755.18 178.10 33 195.40 0

42 01.07.2014 1 933.28 1 769.58 163.70 31 425.82 0

43 01.08.2014 1 933.28 1 773.14 160.14 29 652.68 0

44 01.09.2014 1 933.28 1 782.17 151.11 27 870.51 0

45 01.10.2014 1 933.28 1 795.84 137.44 26 074.67 0

46 *01.11.2014 1 933.28 1 800.41 132.87 24 274.27 0

47 01.12.2014 1 933.28 1 813.57 119.71 22 460.70 0

48 01.01.2015 1 933.28 1 818.82 114.46 20 641.87 0

49 *01.02.2015 1 933.28 1 828.09 105.19 18 813.78 0

50 *01.03.2015 1 933.28 1 846.69 86.59 16 967.10 0

51 01.04.2015 1 933.28 1 846.82 86.46 15 120.28 0

52 01.05.2015 1 933.28 1 858.71 74.57 13 261.57 0

53 01.06.2015 1 933.28 1 865.70 67.58 11 395.86 0

54 01.07.2015 1 933.28 1 877.08 56.20 9 518.78 0

55 *01.08.2015 1 933.28 1 884.77 48.51 7 634.01 0

56 01.09.2015 1 933.28 1 894.38 38.90 5 739.63 0

57 01.10.2015 1 933.28 1 904.98 28.31 3 834.66 0

58 *01.11.2015 1 933.28 1 913.74 19.54 1 920.92 0

59 01.12.2015 1 930.39 1 920.92 9.47 0.00 0

Σ - 115 993.92 100 000.00 15 993.92 100 000.00 0



Differentiated circuit:

1. Calculate the main payment: 100 000/60 = 1666.67

2. We calculate the payments for each month: 1666.67 + 10000 * 6% / 12 = 166.67 + 10000 * 0.005 = 1666.67 + 509.59 = 2176.26

3. We calculate the balance of the debt after the first month: 100 000-1666.67 = 98 333.33

So then, for all 60 months: total loan repayments: 115,247.80

Overpayment amount: 115,247.80-100,000 = 15,247.80


0 *01.01.2011 2 176.26 1 666.67 509.59 98 333.33 0

1 01.02.2011 2 167.76 1 666.67 501.10 96 666.67 0

2 01.03.2011 2 111.60 1 666.67 444.93 95 000.00 0

3 01.04.2011 2 150.78 1 666.67 484.11 93 333.33 0

4 *01.05.2011 2 126.94 1 666.67 460.27 91 666.67 0

5 01.06.2011 2 133.79 1 666.67 467.12 90 000.00 0

6 01.07.2011 2 110.50 1 666.67 443.84 88 333.33 0

7 01.08.2011 2 116.80 1 666.67 450.14 86 666.67 0

8 01.09.2011 2 108.31 1 666.67 441.64 85 000.00 0

9 *01.10.2011 2 085.84 1 666.67 419.18 83 333.33 0

10 01.11.2011 2 091.32 1 666.67 424.66 81 666.67 0

11 01.12.2011 2 069.41 1 666.67 402.74 80 000.00 0

12 *01.01.2012 2 074.30 1 666.67 407.64 78 333.33 0

13 01.02.2012 2 064.75 1 666.67 398.09 76 666.67 0

14 01.03.2012 2 031.15 1 666.67 364.48 75 000.00 0

15 *01.04.2012 2 047.81 1 666.67 381.15 73 333.33 0

16 01.05.2012 2 027.32 1 666.67 360.66 71 666.67 0

17 01.06.2012 2 030.87 1 666.67 364.21 70 000.00 0

18 *01.07.2012 2 010.93 1 666.67 344.26 68 333.33 0

19 01.08.2012 2 013.93 1 666.67 347.27 66 666.67 0

20 *01.09.2012 2 005.46 1 666.67 338.80 65 000.00 0

21 01.10.2012 1 986.34 1 666.67 319.67 63 333.33 0

22 01.11.2012 1 988.52 1 666.67 321.86 61 666.67 0

23 *01.12.2012 1 969.95 1 666.67 303.28 60 000.00 0

24 01.01.2013 1 971.61 1 666.67 304.94 58 333.33 0

25 01.02.2013 1 963.93 1 666.67 297.26 56 666.67 0

26 01.03.2013 1 927.49 1 666.67 260.82 55 000.00 0

27 01.04.2013 1 946.94 1 666.67 280.27 53 333.33 0

28 01.05.2013 1 929.68 1 666.67 263.01 51 666.67 0

29 *01.06.2013 1 929.95 1 666.67 263.29 50 000.00 0

30 01.07.2013 1 913.24 1 666.67 246.58 48 333.33 0

31 01.08.2013 1 912.97 1 666.67 246.30 46 666.67 0

32 *01.09.2013 1 904.47 1 666.67 237.81 45 000.00 0

33 01.10.2013 1 888.58 1 666.67 221.92 43 333.33 0

34 01.11.2013 1 887.49 1 666.67 220.82 41 666.67 0

35 *01.12.2013 1 872.15 1 666.67 205.48 40 000.00 0

36 01.01.2014 1 870.50 1 666.67 203.84 38 333.33 0

37 *01.02.2014 1 862.01 1 666.67 195.34 36 666.67 0

38 *01.03.2014 1 835.43 1 666.67 168.77 35 000.00 0

39 01.04.2014 1 845.02 1 666.67 178.36 33 333.33 0

40 01.05.2014 1 831.05 1 666.67 164.38 31 666.67 0

41 *01.06.2014 1 828.04 1 666.67 161.37 30 000.00 0

42 01.07.2014 1 814.61 1 666.67 147.95 28 333.33 0

43 01.08.2014 1 811.05 1 666.67 144.38 26 666.67 0

44 01.09.2014 1 802.56 1 666.67 135.89 25 000.00 0

45 01.10.2014 1 789.95 1 666.67 123.29 23 333.33 0

46 *01.11.2014 1 785.57 1 666.67 118.90 21 666.67 0

47 01.12.2014 1 773.52 1 666.67 106.85 20 000.00 0

48 01.01.2015 1 768.58 1 666.67 101.92 18 333.33 0

49 *01.02.2015 1 760.09 1 666.67 93.42 16 666.67 0

50 *01.03.2015 1 743.38 1 666.67 76.71 15 000.00 0

51 01.04.2015 1 743.11 1 666.67 76.44 13 333.33 0

52 01.05.2015 1 732.42 1 666.67 65.75 11 666.67 0

53 01.06.2015 1 726.12 1 666.67 59.45 10 000.00 0

54 01.07.2015 1 715.98 1 666.67 49.32 8 333.33 0

55 *01.08.2015 1 709.13 1 666.67 42.47 6 666.67 0

56 01.09.2015 1 700.64 1 666.67 33.97 5 000.00 0

57 01.10.2015 1 691.32 1 666.67 24.66 3 333.33 0

58 *01.11.2015 1 683.65 1 666.67 16.99 1 666.67 0

59 01.12.2015 1 674.89 1 666.67 8.22 0.00 0

Σ - 115 247.80 100 000.00 15 247.80 100 000.00 0